OpenStudy (anonymous):

One of the computer ads Laura is looking at specifies “2 GB Dual Channel DDR2 SDRAM at 800 MHz (max 4 GB).” What does the 2 GB specify?

OpenStudy (anonymous):

It is the Size of the Channel

OpenStudy (rsmith6559):

2 GigaBytes, 2^31 bytes. The 4GB max says that it's a 32 bit system.

OpenStudy (e.mccormick):

G for Gig is sort of like miles. You can have square miles, an area; miles long, a distance; miles per hour, a speed; and so on. It is a specific measure, that as has been pointed out is \(2^{30}\) B (capital) for Bytes is a specific type of measure. If you look at miles in area vs. miles in speed, they are very different types of measure. Thing in bytes are some sort of volume. Right after it tells you it is \(2\times 2^{30}\) bytes, it then says of what: RAM. So, if you look up wjat computer RAM is, you know what you have 2GB of. @rsmith6559 Actually, there have been 64 bit systems with a 4GB max.

OpenStudy (e.mccormick):

@2Chris2 ummm.. what?

OpenStudy (anonymous):

GB = giga byte giga = 1024 * 1 mega mega = 1024 * 1024 so 2 GB = 2 * 1024 * 1024 * 1024 bytes = 2'147'483'648 bytes this is the size of available RAM

OpenStudy (e.mccormick):

And 1024 = \(2^{10}\) so \(1024 \times 1024 \times 1024 = 2^{10} \times 2^{10} \times 2^{10}\). Well, some basic exponent rules apply. \(2^{10} \times 2^{10} \times 2^{10}= 2^{10+10+10} = 2^{30} \)